4. (a)Find the general solution of the partial differential equation.
zxy =(6x^3)y

(b) Find the particular solution for which z(x,0) = 6x^2 , z(1, y) = 6cos y
1
Expert's answer
2020-04-02T13:20:01-0400

a)


"z_{xy}=6x^3y"

"z_x=3x^3y^2+f(x)"

The general solution of the partial differential equation. 

"z={3\over 4}x^4y^2+F(x)+G(y)"

b)

"z(x,0)=6x^2=>F(x)+G(0)=6x^2=>"

"=>F(x)=6x^2, G(0)=0"

"z(1,y)=6\cos{y}=>{3\over 4}y^2+6+G(y)=6\cos{y}=>"

"=>G(y)=-{3 \over 4}y^2+6\cos{y}-6"

"z^*={3\over 4}x^4y^2+6x^2-{3 \over 4}y^2+6\cos{y}-6"